# ESIQA: Perceptual Quality Assessment of Vision-Pro-based Egocentric Spatial Images

### v1

@@ Abstract
With the development of eXtended Reality (XR), head-mounted shooting and display technology have experienced significant advancement and gained considerable attention. Egocentric spatial images and videos are emerging as a compelling form of stereoscopic XR content. Different from traditional 2D images, egocentric spatial images present challenges for perceptual quality assessment due to their special shooting, processing methods, and stereoscopic characteristics. However, the corresponding image quality assessment (IQA) research for egocentric spatial images is still lacking. In this paper, we establish the Egocentric Spatial Images Quality Assessment Database (ESIQAD), the first IQA database dedicated for egocentric spatial images as far as we know. Our ESIQAD includes 500 egocentric spatial images, containing 400 images captured with the Apple Vision Pro and 100 images generated via an iPhone’s “Spatial Camera” app. The corresponding mean opinion scores (MOSs) are collected under three viewing modes, including 2D display, 3D-window display, and 3D-immersive display. Furthermore, based on our database, we conduct a benchmark experiment and evaluate the performance of 22 state-of-the-art IQA models under three different viewing modes. We hope this research can facilitate future IQA research on egocentric spatial images. The database is available at https://github.com/IntMeGroup/ESIQA.
1 Introduction
@@
Egocentric shooting and displaying technology has gained considerable attention recently, which can provide convenient capturing methods and immersive experiences through head-mounted displays (HMDs) [7, 4, 5], such as Apple Vision Pro, Meta Quest 3, and Google Glass, etc. Egocentric spatial images, captured from the first-person perspective, are emerging as an important form of stereoscopic content. As shown in Figure 1, egocentric spatial images are captured with HMDs in real-world scenes and can be transmitted and viewed on various terminal devices. For binocular displays, since a spatial image is composed of a pair of disparity views to simulate human binocular vision, it is possible to directly create a three-dimensional (3D) effect and provide an immersive experience. For conventional flat displays, generally only the left view, i.e., a standard two-dimensional (2D) image, is presented.
Figure 1: Illustration of the capturing and displaying methods for egocentric spatial images.
Egocentric spatial images, taken by a wide variety of individuals from professional photographers to amateurs, exhibit significant differences in visual quality [7, 19]. This quality variation is further compounded by the dynamic and uncontrolled environments when capturing egocentric spatial images, leading to common distortions such as under/over exposure, low visibility, noise, color shift, etc [4]. Moreover, since the captured spatial images can be viewed on any device, the corresponding quality of experience (QoE) may also be different. Therefore, it is important to develop image quality assessment (IQA) metrics of egocentric spatial images to better maintain and optimize QoE on various devices.
@@
Figure 2: Sample egocentric spatial images from our ESIQAD.
@@
Over the past few decades, many studies have focused on the traditional 2D image quality assessment problem, with the establishment of several IQA databases, e.g., TID2013 [23], KonIQ-10K [15], and KADID-10k [16], and the development of many advanced IQA models [21, 18, 35, 31]. However, traditional IQA models are inadequate for assessing the quality of egocentric spatial images due to their unique immersive experiences, which have a distinct impact on QoE. Moreover, spatial images consist of two distinct views (left and right), which precludes the direct application of traditional IQA models. With the diversification of image forms, some studies have focused on 3D IQA (also known as stereoscopic IQA) [22, 10, 8, 36] and VR IQA [28, 3, 37, 6]. However, existing 3D IQA metrics fail to consider the immersive experience of egocentric spatial images, and VR IQA metrics cannot effectively evaluate their stereoscopic effects. These metrics are unsuitable for assessing the quality of egocentric spatial images. As far as we know, there is a lack of IQA studies for egocentric spatial images, which have significant differences compared with traditional 2D images, 3D images, and VR images in terms of formats, characteristics, and applications. Therefore, there is an urgent need for dedicated research on egocentric spatial image quality assessment.
@@
To better understand human visual preferences for egocentric spatial images and facilitate the development of IQA models tailored to these images, we construct the first Egocentric Spatial Images Quality Assessment Database, termed ESIQAD. The database contains 500 egocentric spatial images with diverse scenes and the corresponding collected human perceptual quality ratings on three viewing modes, including 2D, 3D-window, and 3D-immersive. Specifically, 400 spatial images are captured using the Apple Vision Pro, and 100 spatial images are generated by the images taken by the iPhone’s “Spatial Camera” application. Based on the collected egocentric spatial images, we conducted subjective experiments to evaluate the QoE of these images, in which 22 participants were included to provide their perceptual quality ratings under three viewing modes. Based on our ESIQAD, we conduct a benchmark experiment to evaluate the performance of 22 state-of-the-art IQA models under three viewing modes. Our contributions are summarized as follows:
@@
  • We analyze the human preference characteristics for egocentric spatial images under three viewing modes based on our ESIQAD.
  • We conduct a benchmark experiment by evaluating the performance of numerous state-of-the-art IQA models based on our ESIQAD.
@@ 2 Database Construction
To facilitate the egocentric spatial IQA research, we construct a large-scale database, termed ESIQAD, and conduct the corresponding subjective experiment to derive the subjective quality scores in terms of three viewing modes.
Figure 3: Three viewing modes in our subjective experiments.
@@ 2.1 Content Collection
Firstly, we utilized the Apple Vision Pro and the iPhone 12 as the shooting devices to collect spatial images from a variety of scenes, including both indoor and outdoor environments, different weather conditions, and various times of day. These images feature various characteristics including brightness, contrast, colorfulness, spatial details, etc., and have different degradations including noise, blur, etc. The raw images were taken by ten photographers, representing a wide range of user perspectives and preferences. Specifically, we collected 400 spatial images captured with the Apple Vision Pro, and 100 spatial images generated from the images taken by the iPhone, referred to as captured images and generated images, respectively. For the 100 generated images, we used the “Spatial Camera” app to automatically take a 2D image as the left view and then generate the right view to form a spatial image. It should be noted that the 100 generated images have their corresponding images with the same scenes in the 400 images captured by Apple Vision Pro to facilitate the study of human visual preferences. The captured images have a fixed resolution of 2560 × 2560 per view, while the generated images have a resolution of 4032 × 3024 per view. Figure 2 demonstrates sample images from our ESIQAD, where all samples are illustrated in the main view (left view).
@@ 2.2.1 Experiment setup
Based on the collected images, we then conducted a subjective experiment to gather quality scores of human perception on egocentric spatial images across different viewing modes, including 2D, 3D-window, and 3D-immersive, as shown in Figure 3. For the 2D viewing mode, egocentric spatial images were displayed on a 23.8-inch Redmi T238NF monitor with the resolution of 1920 × 1080, driven by a Mac Mini M2 PC. We developed a quality scoring interface using Python Tkinter, which was used to present egocentric spatial images (only the left view) and collect subjective quality ratings. The interface allows users to freely browse the previous and next images, and record the quality scores through a scroll bar. For the 3D-window and 3D-immersive viewing modes, we utilized the Apple Vision Pro device as the head-mounted display, which features the resolution of 3660 × 3200 per eye, the 120°$\degree$ horizontal field of view (FoV), and the 100Hz refresh rate. For the 3D-window mode, egocentric spatial images appeared as small windows within the HMD, allowing users to maintain awareness of the surroundings. For the 3D-immersive mode, egocentric spatial images were shown in a wider immersive FoV, which provides an immersive egocentric experience. In both 3D viewing modes, all images were displayed through the built-in application “Photos” and displayed at their native resolutions to prevent the scaling distortion. Across all three viewing modes, egocentric spatial images were presented in a random order.
@@ 2.2.2 Experiment methodology
We adopted the single stimulus absolute category rating (SSACR) method to collect subjective quality ratings for egocentric spatial images under three viewing modes, with the rating scale ranging from 1 to 10. Under the 2D viewing mode, the viewers were seated at a distance of about 2 feet from the monitor in a laboratory environment with normal indoor illumination, and utilized the scoring interface to assign quality scores. Under the two 3D viewing modes, participants were instructed to wear the Apple Vision Pro and select either the 3D-window or 3D-immersive viewing mode via the “Photos” app. A keyboard was placed in front of the participants to enter the scores. Participants manually entered their scores after viewing each image and proceeded to the next image using gestures.
A total of 22 graduate students (11 males and 11 females) were invited to participate in the subjective experiment. Before the experiment, all participants completed a standard procedure in visual testing, i.e., the Snellen visual acuity test [1]. All participants were confirmed normal or corrected-to-normal vision. For all three viewing modes, the participants were instructed to assess the egocentric spatial images based on both distortion and aesthetic quality, providing ratings with an overall quality score. The whole experiment is split into 2 sessions, each of which includes a subjective quality evaluation process for 750 samples (250 spatial images $\times$ 3 modes). All participants completed both sessions, producing a total of 33,000 quality ratings (22 participants $\times$ 1,500 ratings). After the scoring sessions, each participant was asked to answer a questionnaire regarding their experience. Details of the questionnaire are given in Section 2.3.2.
Figure 4: Distribution of MOSs differences of spatial images across three modes.
@@ 2.3.1 MOS Calculation and Analysis
We perform the same processing methods for the collected scores under three viewing modes respectively. Firstly, we follow the subjective data processing guidelines recommended by ITU [24] for outlier detection and subject rejection. For 2D and 3D-immersive viewing modes, 1 of the 22 participants was deemed as an outlier and excluded, respectively. For the 3D-window viewing mode, none of the 22 participants were excluded. Then, we convert the raw scores from the subjects into normalized Z-scores, which range between 0 and 100. Subsequently, we calculate the average of these Z-scores to derive the mean opinion scores (MOSs), which are formulated as follows:
@@
$\text{MOS}_{j}=\frac{1}{N}\sum_{i=1}^{N}z_{ij}^{\prime},\vspace{-4pt}$ (2)
@@
Figure 5: Distribution of MOSs differences of spatial images across three modes. (a) $\text{MOS}_{\text{3D-window}}-\text{MOS}_{\text{3D-immersive}}$. (b) $\text{MOS}_{\text{3D-immersive}}-\text{MOS}_{\text{2D}}$. (c) $\text{MOS}_{\text{3D-window}}-\text{MOS}_{\text{2D}}$.
In the following analysis, we utilize original MOSs, ranging from 1 to 10, instead of normalized scores to facilitate clearer comparisons across different viewing modes. Figure 4 (a) presents the histograms of the MOS distribution over the three viewing modes, indicating that the perceptual quality scores are widely distributed in the $\left[1,10\right]$ interval in each mode. In addition, Figure 4 (b) and (c) demonstrate the MOS distributions of 400 captured images and 100 generated images, respectively. It can be observed that the MOS distribution for captured images centers around 5 in all three modes, while it centers above 5 for generated images. This is mainly due to the fact that the color of the images captured by the iPhone is generally enhanced to be more comfortable for human vision perception, while the color of the images captured by Apple Vision Pro is generally kept real. Moreover, for captured images, the perceptual quality is higher in 3D-immersive and 3D-window viewing modes compared to 2D viewing mode. However, for generated images, the perceptual quality is better in 2D viewing mode than in 3D viewing modes.
The results shown in Figure 4 (a) indicate that the MOS fitting curves for the two 3D viewing modes are located at the right of the 2D viewing mode, manifesting higher perceived image quality in the 3D environments. To facilitate a more direct comparison of the perceptual quality of the same image across different viewing modes, we calculate the MOS difference for each image between the three modes. The differences are illustrated in the frequency distribution histograms as shown in Figure 5. The results indicate that the perceptual quality of most images in 3D-window and 3D-immersive scenes is comparable but generally superior to that observed on 2D displays.
@@
Figure 6: Distributions of the MOS difference ($\text{MOS}_{\text{generated}}-\text{MOS}_{\text{captured}}$) between 100 matched pairs of captured and generated images under three viewing modes.
@@
To compare the perceptual quality of captured and generated images with identical scenes, we compute the MOS difference between 100 matched pairs of captured and generated images across three viewing modes by subtracting the MOS of generated images and their corresponding captured images. The distribution histogram of the differences is depicted in Figure 6. The results indicate that the perceptual quality of most generated images exceeds that of their corresponding captured images. This improvement mainly results from image enhancement methods applied during the image generation.
Figure 7: A list of the questions included in the post-test questionnaire.

### v2

@@ Abstract
With the development of eXtended Reality (XR), photo capturing and display technology based on head-mounted displays (HMDs) have experienced significant advancements and gained considerable attention. Egocentric spatial images and videos are emerging as a compelling form of stereoscopic XR content. The assessment for the Quality of Experience (QoE) of XR content is important to ensure a high-quality viewing experience. Different from traditional 2D images, egocentric spatial images present challenges for perceptual quality assessment due to their special shooting, processing methods, and stereoscopic characteristics. However, the corresponding image quality assessment (IQA) research for egocentric spatial images is still lacking. In this paper, we establish the Egocentric Spatial Images Quality Assessment Database (ESIQAD), the first IQA database dedicated for egocentric spatial images as far as we know. Our ESIQAD includes 500 egocentric spatial images and the corresponding mean opinion scores (MOSs) under three display modes, including 2D display, 3D-window display, and 3D-immersive display. Based on our ESIQAD, we propose a novel mamba2-based multi-stage feature fusion model, termed ESIQAnet, which predicts the perceptual quality of egocentric spatial images under the three display modes. Specifically, we first extract features from multiple visual state space duality (VSSD) blocks, then apply cross attention to fuse binocular view information and use transposed attention to further refine the features. The multi-stage features are finally concatenated and fed into a quality regression network to predict the quality score. Extensive experimental results demonstrate that the ESIQAnet outperforms 22 state-of-the-art IQA models on the ESIQAD under all three display modes. The database and code are available at https://github.com/IntMeGroup/ESIQA.
keywords:
Egocentric spatial images, quality of experience, image quality assessment, state space model.
[…]
@@
Illustration of the capturing and displaying methods for egocentric spatial images. Captured images, including left and right views, enable standard 2D views on conventional screens and immersive 3D experiences in VR/AR displays. Introduction Egocentric shooting and displaying technology has gained considerable attention recently, which can provide convenient capturing methods and immersive experiences through head-mounted displays (HMDs) [11, 12, 14], such as Apple Vision Pro, Meta Quest 3, and Google Glass, etc. Egocentric spatial images, captured from the first-person perspective, are emerging as an important form of stereoscopic content. As shown in ESIQA: Perceptual Quality Assessment of Vision-Pro-based Egocentric Spatial Images, egocentric spatial images are captured with HMDs in real-world scenes and can be transmitted and viewed on various terminal devices. For binocular displays, since a spatial image is composed of a pair of disparity views to simulate human binocular vision, it is possible to directly create a three-dimensional (3D) effect and provide an immersive experience in HMDs. For conventional flat displays, generally only the left view, i.e., a standard two-dimensional (2D) image, is presented.
@@
Figure 1: Sample egocentric spatial images from our ESIQAD, where all samples are illustrated in their left view.
Egocentric spatial images, taken by a wide variety of individuals from professional photographers to amateurs, exhibit significant differences in visual quality [14, 31]. This quality variation is further compounded by the dynamic and uncontrolled environments when capturing egocentric spatial images [11], leading to common distortions such as under/over exposure, low visibility, noise, color shift, etc. Moreover, since the captured egocentric spatial images can be viewed on any device, the corresponding Quality of Experience (QoE) may also be different. In order to ensure the QoE of end-users, the service providers need to monitor the quality of egocentric spatial images in the entire media stream, including uploading, compressing, post-processing, and transmission. With thousands of egocentric spatial images being captured, uploaded and viewed every day, it is crucial to develop specialized image quality assessment (IQA) metrics for egocentric spatial images to better maintain and optimize the QoE.
@@
Over the past few decades, many studies have focused on the traditional 2D image quality assessment problem, with the establishment of several IQA databases, e.g., TID2013 [39], KonIQ-10K [28], and KADID-10k [29], and the development of many advanced IQA models [33, 35, 49, 54]. However, traditional IQA models are inadequate for assessing the quality of egocentric spatial images due to their unique immersive experiences, which have a distinct impact on QoE. Moreover, spatial images consist of two distinct views (left and right), which precludes the direct application of traditional IQA models. With the diversification of image forms, some studies have focused on 3D IQA (also known as stereoscopic IQA) [15, 17, 36, 55] and VR IQA [10, 13, 45, 56, 26]. However, existing 3D IQA metrics are mainly designed for traditional stereoscopic images, which rely on dual-eye views to create depth perception. These metrics take 3D effects into account but overlook the immersive experience of egocentric spatial images (which generally requires a larger field of view (FoV) compared to classical 2D/3D small display) and the first-person perspective in dynamic environments. Meanwhile, VR IQA metrics are typically applied to VR images such as omnidirectional images in immersive environments. However, these metrics generally ignore the stereoscopic depth effects and are not designed for the first-person egocentric perspective. As far as we know, there is a lack of IQA studies for egocentric spatial images, which have significant differences compared with traditional 2D images, 3D images, and VR images in terms of formats, characteristics, and applications. Therefore, there is an urgent need for dedicated research on egocentric spatial image quality assessment from both subjective and objective perspectives.
@@
To better understand human visual preferences for egocentric spatial images and facilitate the development of IQA models tailored to these images, we construct the first Egocentric Spatial Images Quality Assessment Database, termed ESIQAD. The database contains 500 egocentric spatial images with diverse scenes and the corresponding collected human perceptual quality ratings under three display modes, including 2D display, 3D-window display, and 3D-immersive display. Specifically, 400 egocentric spatial images are captured using the Apple Vision Pro, and 100 egocentric spatial images are taken using the iPhone and synthesized by the iPhone’s “Spatial Camera" application. Based on the collected images, we conducted subjective experiments to evaluate the QoE of them, in which 22 participants were included to provide their perceptual quality ratings under three display modes. Based on our ESIQAD, we introduce a novel mamba2-based multi-stage feature fusion model, which leverages the effectiveness of the state space models (SSM) to predict the perceptual quality of egocentric spatial images across three display modes, termed ESIQAnet. We first input the stereo image pairs into a patch embedding layer. The model then extracts features using visual state space duality (VSSD) blocks in the first three stages and multi-head self attention (MSA) blocks in the final stage. For the extracted multi-scale features, we apply cross attention and transposed attention to fuse and enhance visual information from both views. The outputs are then averaged and concatenated to form a comprehensive feature representation, which is fed into a quality regression network to predict the quality score. We validate the performance of the proposed model and 22 state-of-the-art benchmark IQA models on the ESIQAD. Extensive experimental results demonstrate that our model outperforms the benchmark IQA models and has good generalization performance for different display modes. Our contributions are summarized as follows:
@@
  • We analyze the human preference characteristics for egocentric spatial images under three display modes based on our ESIQAD.
  • We conduct a benchmark experiment by evaluating the performance of numerous state-of-the-art IQA models.
  • We propose ESIQAnet, which is built based on mamba2 architecture to perform egocentric spatial image quality assessment. Our ESIQAnet outperforms other benchmark models across all three display modes.
@@ 1 Database Construction
To facilitate the egocentric spatial IQA research, we construct a large-scale database, termed ESIQAD, and conduct the corresponding subjective experiment to derive the subjective quality scores in terms of three display modes.
@@ 1.1 Content Collection
Firstly, we utilized the Apple Vision Pro and the iPhone 12 as the shooting devices to collect egocentric spatial images from a variety of scenes, including both indoor and outdoor environments, different weather conditions, and various times of day. These images feature various characteristics including brightness, contrast, colorfulness, spatial details, etc., and have different degradations including noise, blur, etc. The raw images were taken by ten photographers, representing a wide range of user perspectives and preferences. Specifically, we collected 400 egocentric spatial images captured with the Apple Vision Pro, and 100 egocentric spatial images synthesized from the images taken by the iPhone, referred to as captured images and synthesized images, respectively. For the 100 synthesized images, we used the wide and ultra-wide cameras of iPhone 12 to simultaneously capture two images, serving as the left and right views for synthesizing a spatial image. When the iPhone is held horizontally, the two cameras can mimic human eye parallax. It should be noted that the 100 synthesized images have their corresponding images with the same scenes in the 400 images captured by Apple Vision Pro to facilitate the study of human visual preferences. The captured images have a fixed resolution of 2560 × 2560 per view, while the synthesized images have a fixed resolution of 4032 × 3024 per view. Fig. 1 demonstrates sample images from our ESIQAD, where all samples are illustrated in the main view (left view).
@@ 1.2 Subjective Experiment
Figure 2: Three display modes in our subjective experiments.
@@ 1.2.1 Experiment setup
Based on the collected images, we then conducted a subjective experiment to gather quality scores of human perception on egocentric spatial images across different display modes, including 2D, 3D-window, and 3D-immersive, as shown in Fig. 2. For the 2D display mode, egocentric spatial images were displayed on an iMac monitor with a resolution of 4K, driven by a Mac Mini M2 PC. We developed a quality scoring interface using Python Tkinter, which was used to present egocentric spatial images (only the left view) and collect subjective quality ratings. The interface allows users to freely browse the previous and next images, and record the quality scores through a scroll bar. For the 3D-window and 3D-immersive display modes, we utilized the Apple Vision Pro device as the HMD, which features a resolution of 3660 × 3200 per eye, a 120°$\degree$ horizontal FoV, and a 100Hz refresh rate. For the 3D-window mode, egocentric spatial images appeared as small windows within the HMD, allowing users to maintain awareness of the surroundings. For the 3D-immersive mode, egocentric spatial images were shown in a wider immersive FoV, which provides an immersive egocentric experience. In both 3D display modes, all images were displayed through the built-in application “Photos” and displayed at their native resolutions to prevent the scaling distortion. Across all three display modes, egocentric spatial images were presented in a random order.
@@ 1.2.2 Experiment methodology
We adopted the single stimulus absolute category rating (SSACR) method to collect subjective quality ratings for egocentric spatial images under three display modes, with the rating scale ranging from 1 to 10. Under the 2D display, the viewers were seated at a distance of about 2 feet from the monitor in a laboratory environment with normal indoor illumination and utilized the scoring interface to assign quality scores. Under the two 3D displays, participants were instructed to wear the Apple Vision Pro and select either the 3D-window or 3D-immersive display mode via the “Photos" app. A keyboard was placed in front of the participants to enter the scores. Participants manually entered their scores after viewing each image and proceeded to the next image using gestures.
According to the recommendation from ITU-R BT.500-13 [1] and BT.2021 [2], we invited a total of 22 graduate students to participate in the subjective experiment. Before the experiment, all participants completed a survey to provide the participant demographics. Among all participants (11 males and 11 females, age range 22–30, average age 25), 14 indicated that they spend no time in HMDs per week, while 6 indicated that they spend between 0.5-3 hours, and 2 indicated they spend > 3 hours. All participants were required to complete a standard procedure in visual testing, i.e., the Snellen visual acuity test [3], where 8 of them were confirmed to have normal vision and 14 had corrected-to-normal vision (9 wore contact lenses, 2 wore glasses, 2 wore neither, and 1 had undergone refractive surgery). For all three display modes, the participants were instructed to assess the egocentric spatial images based on both distortion and aesthetic quality, providing ratings with an overall quality score. The whole experiment was split into 2 sessions, each of which included a subjective quality evaluation process for 750 samples (250 spatial images $\times$ 3 display modes). All participants completed both sessions, producing a total of 33,000 quality ratings (22 participants $\times$ 1,500 ratings). After the scoring sessions, each participant was asked to answer a post-study questionnaire regarding their experience. Details of the questionnaire are given in Section 1.3.2.
Figure 3: MOS discriminability and mean CI evolution with participants’ number in ESIQAD.
Figure 4: Distribution of MOSs of egocentric spatial images across three display modes.
@@ 1.3.1 MOS Calculation and Analysis
We performed the same processing methods for the collected scores under three display modes respectively. Firstly, we followed the subjective data processing guidelines recommended by ITU [40] for outlier detection and subject rejection. For 2D and 3D-immersive display modes, 1 of the 22 participants was deemed as an outlier and excluded, respectively. For the 3D-window display mode, none of the 22 participants were excluded. Then, we convert the raw scores from the subjects into normalized Z-scores, which range between 0 and 100. Subsequently, we calculate the average of these Z-scores to derive the mean opinion scores (MOSs), which are formulated as follows:
@@
$\text{MOS}_{j}=\frac{1}{N}\sum_{i=1}^{N}z_{ij}^{\prime},$ (2)
@@
To investigate the reliability of MOSs, we calculated the discriminability and mean Confidence Interval (CI) metrics of the ESIQAD with increasing assessor numbers [37, 38]. For the discriminability metric, we applied the two-sample Wilcoxon test on all possible pairs of MOSs in the ESIQAD to test the ratio of significantly different ones according to [38]. The mean CI was calculated by averaging the standard deviations of scores for each sample (scaled by a Z-score for a 95% confidence level) across all samples, which measures the average uncertainty around the MOSs within the database. Fig. 3 illustrates the trend of the discriminability and mean CI metrics as the number of participants changes, showing that 22 participants can yield relatively reliable MOSs.
@@
Figure 5: Distribution of MOS difference of spatial images across three modes. (a) $\text{MOS}_{\text{3D-window}}-\text{MOS}_{\text{3D-immersive}}$. (b) $\text{MOS}_{\text{3D-immersive}}-\text{MOS}_{\text{2D}}$. (c) $\text{MOS}_{\text{3D-window}}-\text{MOS}_{\text{2D}}$.
@@
Figure 6: Distributions of the MOS difference ($\text{MOS}_{\text{synthesized}}-\text{MOS}_{\text{captured}}$) between 100 matched pairs of captured and synthesized images under three display modes.
Fig. 4(a) presents the histograms of the MOS distribution over the three viewing modes, indicating that the perceptual quality scores are widely distributed in the $\left[1,10\right]$ interval in each mode. In addition, Fig. 4(b) and Fig. 4(c) demonstrate the MOS distributions of 400 captured images and 100 synthesized images, respectively. It can be observed that the MOS distribution for captured images centers around 5 in all three modes, while it centers above 5 for synthesized images. This is mainly because the image captured by the iPhone has a higher fixed resolution and their color is generally enhanced to be more comfortable for human vision perception. Moreover, for captured images, the perceptual quality is higher in 3D-immersive and 3D-window viewing modes compared to the 2D viewing mode. However, for synthesized images, the perceptual quality is better in 2D viewing mode than in 3D viewing modes.
The results shown in Fig. 4(a) indicate that the MOS fitting curves for the two 3D viewing modes are located at the right of the 2D viewing mode, manifesting higher perceived image quality in the 3D environments. To facilitate a more direct comparison of the perceptual quality of the same image across different viewing modes, we calculate the MOS difference for each image between the three modes. The differences are illustrated in the frequency distribution histograms as shown in Fig. 6. The results indicate that the perceptual quality of most images in 3D-window and 3D-immersive scenes is comparable but generally superior to that observed on 2D displays. This may be attributed to the immersive and stereoscopic nature of 3D viewing modes, which offer better senses with depth and realism. In 3D environments, viewers experience enhanced spatial details and a more real scene representation, leading to an overall improvement in perceived image quality compared to 2D displays.
@@
To compare the perceptual quality of captured and synthesized images with identical scenes, we compute the MOS difference between 100 matched pairs of captured and synthesized images across three display modes by subtracting the MOS of synthesized images and their corresponding captured images. The distribution histogram of the differences is depicted in Fig. 6. The results indicate that the perceptual quality of most synthesized images exceeds that of their corresponding captured images. This improvement mainly results from the higher resolution and more accurate color reproduction in synthesized images taken by phones.
Figure 7: Questions included in the post-test questionnaire.
Figure 9: Illustration of (a) the overall architecture of the ESIQAnet, (b) the structure of visual state space duality (VSSD) block, and (c) the structure of transposed attention block.